# Co-Paced Learning Strategy Based on Confidence for Flying Bird Object Detection Model Training

### v1

@@ Abstract
To mitigate the adverse effects of hard samples on the training of the Flying Bird Object Detection (FBOD) model for surveillance videos, we propose a Co-Paced Learning Based on Confidence (CPL-BC) strategy and apply this strategy to the training process of the FBOD model. This strategy involves maintaining two models with identical structures but different initial parameter configurations, which collaborate with each other to select easy samples with prediction confidence exceeding a set threshold for training. As training progresses, the strategy gradually lowers the threshold, allowing more samples to participate, enhancing the model’s ability to recognize objects from easy to hard. Before applying the CPL-BC strategy to train the FBOD models, we initially trained the two FBOD models to equip them with the capability to assess the difficulty level of flying bird object samples. Experimental results on two different datasets of flying bird objects in surveillance videos demonstrate that, compared to other model learning strategies, CPL-BC significantly improves detection accuracy, verifying the effectiveness and advancement of this method.
@@
[inst1]organization=School of Information Science and Technology, Southwest Jiaotong University,city=Chengdu, postcode=610000, country=China
@@ 1 Introduction
Detection of flying bird objects has essential application value in many fields [1, 2, 3]. At present, the method of detecting flying bird objects by radar is widely used [4, 5], but the radar equipment has some disadvantages, such as large size, expensive, and poor visual effect. With the development of computer vision, deep learning, and other technologies, there is increasing research on using cameras to detect various objects [6, 7, 8].
@@
We are working on how to use surveillance cameras to detect flying bird objects [9, 10, 11, 12]. However, the flying bird objects captured by surveillance cameras present different identification difficulties due to their different sizes or different similarities to the background. As shown in Figure 1, these flying bird objects have various levels of difficulty to identify. In the training process of the model, if easy samples and hard samples are not distinguished, the model may be affected by the noise in hard samples, thus reducing the detection accuracy [12].
Figure 1: Identifying the flying bird object in the surveillance video has different degrees of difficulty.
@@
To solve the above problem, an intuitive approach is to train the model using only easy samples and discard the hard ones (this approach is similar to filtering out the noisier samples when processing data with noise labels [13, 14, 15, 16]). In this way, it can avoid the interference of fuzzy, hard samples with the model training process to a certain extent, thus reducing the possibility of false detection. However, because hard samples are not included in the training process, the model may perform poorly when faced with these hard samples. To solve this contradiction, a new method, Self-Paced Learning strategy with Easy Sample Prior Based on Confidence (SPL-ESP-BC), was proposed in the literature [12]. The core idea of this strategy is to start learning from easy samples and then gradually transition to hard samples to gradually improve the model’s recognition ability of the object and effectively reduce the adverse impact of hard samples on model training. Compared with the loss-based Self-Paced Learning (SPL) strategy [17, 18, 19, 20, 21, 22], the SPL-ESP-BC [12] uses the prediction confidence of samples to evaluate the difficulty degree of samples. In the One-Category object detection task (the flying bird object detection belongs to the One-Category object detection), using confidence as a criterion has greater advantages than the loss of use. However, it should be noted that in the process of SPL, bias accumulation may occur in sample selection. Specifically, the model may mistakenly treat some hard samples as easy samples or the easy samples as hard samples, resulting in selection bias. Since the SPL strategy relies on the model itself to evaluate the difficulty of the sample, the sample selection bias may gradually accumulate with the training process and eventually affect the model’s performance.
@@
In response to the challenge of noise labeling, literature [23] proposes a novel and effective learning paradigm called “Co-teaching”. This paradigm trains two deep neural networks at the same time. After each iteration completes the forward inference, each network will select a batch of data with sufficiently small losses according to their calculated loss value and consider these data to be more accurate (because these data have less loss when passing through the network, meaning that the network’s prediction results for these data are more consistent with the given label). These batches of data, which are considered “clean”, are then passed on to the peer network for it to update its parameters. Since deep networks tend to preferentially learn “clean” data and become robust [24], when data containing noise is mistakenly considered “clean” and passed to a peer network, the peer network can screen out these noisy data to some extent by its robustness, effectively preventing the accumulation of this type of selection bias. In this mutual screening and transmission mechanism, the two networks can work together to reduce the errors introduced by noise labels.
@@
Inspired by Co-teaching and combined with the principle of Self-Paced Learning strategy Based on Confidence (SPL-BC), In this paper, a Co-Paced Learning stragegy Based on Confidence (CPL-BC) is proposed to optimize the training process of Flying Bird Object Detection (FBOD) model for surveillance video. The core of this strategy is to maintain two object detection models with the same structure but different initial parameter configurations. In each training iteration, the two models independently evaluate the confidence of the sample set and select those whose confidence exceeds the currently set threshold as “easy samples”. These selected easy samples are then passed to each other’s networks for parameter updates. As the training process progresses and the number of training rounds increases, the confidence threshold is gradually lowered to allow more samples (which gradually become hard) to participate in the training. This process starts with easy samples and progressively transitions to hard samples, which can enhance the model’s ability to recognize object features and reduce the adverse effects of hard samples on the model training process. By alternately assessing the difficulty level of samples using two models, this strategy can, to some extent, prevent the accumulation of sample selection bias.
@@
Before applying the CPL-BC strategy to the training of the FBOD model, we first independently and randomly initialize two FBOD models with the same structure to ensure that they have different initial parameter configurations. Then, the two models are trained preliminarily so that each can evaluate the difficulty degree of samples. Then, the CPL-BC strategy is applied to the subsequent training of the two models.
@@
    The CPL-BC model training strategy for the One-Category object detection model is proposed by integrating the concept of Co-teaching into the SPL-BC strategy. The strategy involves training two neural network models simultaneously, each of which independently evaluates the confidence of the sample in each training iteration. They select those samples whose confidence exceeds the currently set threshold as “easy samples” and pass those samples to the peer network for parameter updates. With the training process’s deepening and increasing training rounds, the confidence threshold is gradually lowered to allow more samples to participate in the training. This strategy is designed to avoid the accumulation of selection bias when relying on a single model for evaluation.
@@
    Applying the CPL-BC strategy to the training process of the FBOD model. First, two FBOD models with the same structure are initialized independently and randomly, ensuring they have different initial parameter configurations to exhibit different initial learning capabilities. Then, the two models are preliminarily trained to evaluate the difficulty level of the samples. Then, in the subsequent training stage, the CPL-BC strategy is adopted to improve the detection performance of the FBOD model further. Specifically, the CPL-BC strategy makes the model gradually pay attention to the more difficult-to-identify samples in training by alternately selecting samples and gradually adjusting the confidence threshold, thus improving its overall detection performance.
@@
2 Related Works
The main work of this paper is to combine the Co-teaching method with the SPL-BC model training strategy innovatively and apply this fusion strategy to the training process of the FBOD model for surveillance video. To facilitate the introduction of subsequent content, in this section, we briefly review the SPL-BC method and the surveillance video flying bird object detection method FBOD-SV [10].
@@ 2.1 The SPL-BC Algorithm
where $\lambda$ is the age parameter that controls the learning speed. $\text{L}\left(\text{y}_{i},f\left(\textbf{x}_{i};\textbf{w}\right)\right)$ represents the loss between the predicted result and the label. $_{i}\in\textbf{v}$ v $\left(\textbf{v}=[v_{1},v_{2},\cdots,v_{n}]\right)$ is the sample weight, weighted to the loss of the $i^{th}$ sample, which is used to control whether the sample participates in training or the degree of participation. $g\left(v_{i},\lambda\right)$ is the Self-Paced Regularizer, and the determination of the sample weight depends on it.
@@
Hard [17] $-\lambda\sum\limits_{i=1}^{n}v_{i}$$\begin{cases}1,&l_{i}<\lambda,\\ 0,&{\text{ Otherwise}},\end{cases}$
@@
In the process of SPL-BC, the parameter $\xi$ gradually decreases as the number of training iterations increases. This implies that samples initially considered hard due to their lower prediction as training progresses confidence will be more easily included in the training process due to the reduction in the $\xi$ threshold. Specifically, $\xi$ can be viewed as a dynamic threshold parameter closely related to the prediction confidence of samples, determining which hard samples (i.e., samples with lower prediction confidence) will be included in the current training round. This adjustment strategy, where $\xi$ gradually decreases with increasing training iterations, is referred to as the Training Schedule strategy.
@@ 2.2 The FBOD-SV Method
The FBOD-SV [10] is designed to detect flying birds in surveillance videos. Its input consists of consecutive frames of video images, and its output is the position information of flying birds on these video images, represented specifically as bounding boxes. The implementation workflow of FBOD-SV is as follows: Firstly, a processing unit focused on fusing multi-frame video image information, known as the Correlation Attention Feature Aggregation unit (Co-Attention-FA), is utilized to integrate this information, with special attention given to flying bird object in the video. Subsequently, a network structure that first down-samples and then up-samples is employed to extract features of the flying bird objects. These features are then passed to a large feature map, which is used for subsequent flying bird object prediction. Finally, the model outputs two branches: one branch is responsible for predicting whether feature points (i.e., anchor points) in the feature map belong to a flying bird object and providing the corresponding confidence score (the confidence prediction branch); the other branch is responsible for providing the specific location information of the flying bird object relative to the feature points, i.e., the regression results of the bounding box (the bounding box regression branch). The FBOD-SV employs a multi-task loss function to train the FBOD model. This multi-task loss function consists of confidence loss and location regression loss. Specifically, during the training process, the loss value of each anchor point is obtained by calculating the weighted sum of its confidence loss and location regression loss. The loss function formula is as follows,
@@
Total Loss $=\frac{1}{\text{N}}\sum{\text{L}\left(\text{A}_{i}\right)}$
@@
Total $\text{Total Loss}=\frac{1}{\text{N}}\left(\text{L}\left(\text{A}_{neg}\right)+\text{L}\left(\text{A}_{F_{1}}\right)+...+\text{L}\left(\text{A}_{F_{n}}\right)\right),$ (10)
@@ 3 The CPL-BC Method
The CPL-BC strategy involves simultaneously maintaining two models that work together to pick easy samples for each other whose prediction confidence exceeds a certain threshold and train them. As the training process deepens, the method adopts certain strategies to gradually reduce this threshold in order to allow more samples to participate in the training. Specifically, the CPL-BC strategy consists of the following five steps.
@@
Firstly, define a confidence-based Minimizer Function $\sigma^{\prime}$, which is used to determine the optimal weights for training samples. Next, establish a Training Schedule method $\Psi$ that will be employed to dynamically adjust the criteria for evaluating the easiness or hardness of samples during the training process. Additionally, set an initial confidence threshold-related parameter $\xi$ as the starting point. Subsequently, initialize two models, $f$ and $g$, with identical structures using two different sets of weight parameters,$\textbf{w}_{f}$ and $\textbf{w}_{g}$, respectively.
@@ 4 Applying the CPL-BC strategy to the training process of the FBOD model
Figure 2 shows the training process of the FBOD model for surveillance video by applying the proposed CPL-BC strategy in this paper. The process consists of two main stages: the Model Prior stage and the CPL-BC training stage. Specifically, in the Model Prior stage, all samples (without distinguishing the difficulty degree of samples) are used to train two FBOD models (the weights of both models are randomly initialized) so that they initially can determine the difficulty degree of samples [as shown in Fig. Figure 2(a)]. In the CPL-BC learning stage, all samples are used again, and the CPL-BC strategy is adopted for further training of the model (at this time, the model weights are initialized using the fixed model weights after the Model Prior stage)[as shown in Fig. Figure 2(b)]. In this stage, the predictive confidence of one model for a sample is used to determine the optimal sample weight of the Weighted Loss Function of the other model when processing the sample, thus controlling whether the sample participates in the training (or the degree of participation).
@@ 4.1 The Weighted Loss Function
Total Loss $=\frac{1}{\text{N}}\left(\text{L}\left(\text{A}_{neg}\right)+v_{i}\text{L}\left(\text{A}_{F_{1}}\right)+...+v_{n}\text{L}\left(\text{A}_{F_{n}}\right)\right)$
$=\frac{1}{\text{N}}\left(\text{L}\left(\text{A}_{neg}\right)+\boldsymbol{\vec{v}}\left(\text{L}\left(\text{A}_{F_{1}}\right)...\ \text{L}\left(\text{A}_{F_{n}}\right)\right)^{T}\right),$ (16)
@@ 4.3 The CPL-BC strategy for the FBOD model training
Algorithm 1 presents the pseudocode for applying the CPL-BC strategy to the training process of the FBOD model. Initially, two identical-structure FBOD models are pre-trained using the entire dataset (model priors). During this pre-training phase, a conventional training strategy is employed, which does not differentiate between the difficulty levels of samples nor weigh their losses. Training ceases before the models overfit to hard samples, typically implying a relatively low number of training iterations, to ensure the models initially possess the ability to distinguish between easy and hard samples. Subsequently, the CPL-BC strategy is adopted to train these two models further. This strategy aims to enable the models to progressively learn features of flying bird objects, starting from easy samples and progressing to hard ones. In the process of CPL-BC, the two models assess the difficulty level of samples for each other, meaning that one model evaluates the ease or difficulty of a sample for the other based on its prediction confidence. This approach can effectively mitigate the issue of cumulative bias that may arise when a single model selects samples.

### v2

# Co-Paced Learning Strategy Based on Confidence for the Flying Bird Object Detection Model Training

@@ Abstract
The flying bird objects captured by surveillance cameras exhibit varying levels of recognition difficulty due to factors such as their varying sizes or degrees of similarity to the background. To alleviate the negative impact of hard samples on training the Flying Bird Object Detection (FBOD) model for surveillance videos, we propose the Co-Paced Learning strategy Based on Confidence (CPL-BC) and apply it to the training process of the FBOD model. This strategy involves maintaining two models with identical structures but different initial parameter configurations that collaborate with each other to select easy samples for training, where the prediction confidence exceeds a set threshold. As training progresses, the strategy gradually lowers the threshold, thereby gradually enhancing the model’s ability to recognize objects, from easier to more hard ones. Prior to applying CPL-BC, we pre-trained the two FBOD models to equip them with the capability to assess the difficulty of flying bird object samples. Experimental results on two different datasets of flying bird objects in surveillance videos demonstrate that, compared to other model learning strategies, CPL-BC significantly improves detection accuracy, thereby verifying the method’s effectiveness and advancement.
@@
[…]
@@ 1 Introduction
The detection of flying bird objects holds significant application value across various fields [1, 2, 3]. Currently, the radar-based method for detecting flying bird objects is widely used [4, 5], but radar equipment suffers from disadvantages such as large size, high cost, and limited visual effectiveness. With the advancement of computer vision, deep learning, and other technologies, there is a growing amount of research focused on using cameras to detect various objects [6].
@@
We are working on how to use surveillance cameras to detect flying bird objects [7, 8, 9, 10]. However, the flying bird objects captured by surveillance cameras pose different levels of identification difficulty due to their varying sizes or degrees of similarity to the background. During model training, the model may be adversely affected by the noise inherent in hard samples (e.g., those illustrated in Figure 1), thereby reducing detection accuracy [10].
Figure 1: Examples of hard samples of flying bird objects (red boxes indicate flying bird objects)
@@
To address the aforementioned problem, an intuitive approach would be to train the model using only easy samples and exclude the hard ones (similar to filtering out noisy samples when dealing with data labeled with noise [11, 12, 13, 14]). By doing so, it can, to a certain extent, avoid the interference of ambiguous, hard samples during the model training process, thereby reducing the likelihood of false detections. However, excluding hard samples from the training process may result in poor performance when the model encounters these hard samples. To resolve this contradiction, a new method called the Self-Paced Learning strategy with Easy Sample Prior Based on Confidence (SPL-ESP-BC) was proposed in the literature [10]. The core idea of this strategy is to initiate learning from easy samples and gradually transition to hard samples, improving the model’s recognition ability of the object step by step and effectively mitigating the adverse impact of hard samples on model training. Compared to the loss-based Self-Paced Learning (SPL) strategy [15, 16, 17, 18, 19, 20], SPL-ESP-BC [10] utilizes the prediction confidence of samples to assess their difficulty level. In the context of One-Category object detection tasks (where flying bird object detection falls), using confidence as a criterion offers greater advantages than using loss. However, it should be noted that during the SPL process, bias accumulation may arise in sample selection. Specifically, the model may mistakenly categorize some hard samples as easy ones or vice versa, leading to selection bias. Since the SPL strategy relies on the model itself to evaluate the sample difficulty, the sample selection bias may gradually accumulate during the training process and ultimately affect the model’s performance.
@@
In response to the challenge of noisy labeling, literature [21] proposes a novel and effective learning paradigm called “Co-teaching”. This paradigm involves training two deep neural networks simultaneously. After each iteration completes its forward inference, each network selects a batch of data with sufficiently small losses based on its calculated loss value, considering these data to be more accurate (as they exhibit less loss when processed through the network, indicating that the network’s predictions for these data are more consistent with the given labels). These batches of data, deemed “clean”, are then passed to the peer network for parameter updates. Since deep networks have a tendency to prioritize learning from “clean” data and develop robustness [22], when data containing noise is mistakenly identified as “clean” and passed to a peer network, the peer network can, to a certain extent, filter out these noisy data due to its robustness, thereby effectively preventing the accumulation of such selection bias. Through this mutual screening and transmission mechanism, the two networks collaborate to reduce the errors introduced by noisy labels.
@@
Inspired by Co-teaching and combined with the principle of Self-Paced Learning strategy Based on Confidence (SPL-BC), this paper proposes a Co-Paced Learning strategy Based on Confidence (CPL-BC) to optimize the training of the Flying Bird Object Detection (FBOD) model for surveillance video. The core of this strategy is maintaining two object detection models with the same structure but different initial parameter configurations. In each training iteration, the models independently evaluate the confidence of the sample set and select those whose confidence exceeds the currently set threshold as “easy samples”. These selected easy samples are then exchanged between the models for parameter updates. As the training progresses and the number of rounds increases, the confidence threshold is gradually lowered to include more samples, gradually transitioning from easy to hard. This process starts with easy samples and progressively incorporates harder ones, enhancing the model’s ability to recognize object features and mitigating the adverse effects of hard samples on training. By alternately assessing sample difficulty using two models, this strategy can, to some extent, prevent the accumulation of sample selection bias.
@@
Before applying the CPL-BC strategy to train the FBOD model, we first independently and randomly initialize two FBOD models with the same structure to ensure they have distinct initial parameter configurations. Then, both models undergo preliminary training to equip them with the ability to assess the difficulty of samples. Subsequently, the CPL-BC strategy is applied during their further training.
@@
    The CPL-BC training strategy for the One-Category object detection model is proposed by integrating the concept of Co-teaching into the SPL-BC strategy. It involves simultaneously training two neural network models, each independently evaluating sample confidence in each iteration. They select “easy samples” with confidence exceeding the current threshold and pass them to the peer network for updates. As training progresses and rounds increase, the threshold gradually lowers, allowing more samples to train. This avoids selection bias from single-model evaluation.
@@
    The CPL-BC strategy is applied to the training process of the FBOD model. Firstly, two FBOD models with identical structures are independently and randomly initialized to ensure they possess different initial parameter configurations. Subsequently, these two models undergo preliminary training to equip them with the capability to assess the difficulty levels of samples. Then, in the subsequent training phase, the CPL-BC strategy is adopted to further enhance the detection performance of the FBOD model.
@@
2 Preliminaries
The main focus of this paper is to innovatively combine the Co-teaching method with the SPL-BC model training strategy, and apply this integrated strategy to the training process of the FBOD model for surveillance video. To pave the way for the subsequent content, in this section, we briefly review the SPL-BC method and introduce the FBOD-SV method [8], which is specifically designed for flying bird object detection in surveillance videos.
@@ 2.1 The SPL-BC Algorithm
where $\lambda$ is the age parameter that controls the learning speed. $\text{L}\left(\text{y}_{i},f\left(\textbf{x}_{i};\textbf{w}\right)\right)$ represents the loss between the predicted result and the corresponding label. $v_{i}\in\textbf{v}$ $\left(\textbf{v}=[v_{1},v_{2},\cdots,v_{n}]\right)$ is the sample weight, weighted to the loss of the $i^{th}$ sample, which is used to control whether the sample participates in training or the degree of participation. $g\left(v_{i},\lambda\right)$ is the Self-Paced Regularizer, and the determination of the sample weight depends on it.
@@
Hard [15] $\lambda\sum\limits_{i=1}^{n}v_{i}$ $\begin{cases}1,&l_{i}<\lambda,\\ 0,&{\text{ Otherwise}},\end{cases}$
@@
Polynomial[23] $\begin{array}[c]{@{}c@{}}\lambda\left(\frac{1}{t}\sum\limits_{i=1}^{N}v_{i}^{t}-\sum\limits_{i=1}^{N}v_{i}\right)\\ \lambda>0,t>1\end{array}$\$} {(1-liλ)1t-1,li<λ,0, Otherwise,
@@
In the process of SPL-BC, the parameter $\xi$ gradually decreases as the number of training iterations increases. This implies that samples initially considered hard, due to their lower prediction confidence, will become more likely to be included in the training process as the threshold $\xi$ decreases. Specifically, $\xi$ can be viewed as a dynamic threshold parameter closely related to the prediction confidence of samples, determining which hard samples (i.e., samples with lower prediction confidence) are included in the current training round. This adjustment strategy, where $\xi$ gradually decreases with the increasing number of training iterations, is referred to as the Training Schedule strategy.
@@ 2.2 The FBOD-SV Method
The FBOD-SV [8] is designed to detect flying birds in surveillance videos. Its input comprises consecutive frames of video images, and its output is the position information of flying birds on these images, specifically represented as bounding boxes. The implementation workflow of FBOD-SV is as follows: Firstly, a processing unit focused on fusing multi-frame video image information, known as the Correlation Attention Feature Aggregation unit (Co-Attention-FA), is utilized to integrate this information, with special attention given to the flying bird objects in the video. Subsequently, a network structure that first down-samples and then up-samples is employed to extract features of the flying bird objects. These features are then passed to a large feature map for subsequent prediction of flying bird objects. Finally, the model outputs two branches: one branch is responsible for predicting whether feature points (i.e., anchor points) in the feature map belong to a flying bird object and providing the corresponding confidence score (the confidence prediction branch); the other branch is responsible for providing the specific location information of the flying bird object relative to the feature points, i.e., the regression results of the bounding box (the bounding box regression branch).
The FBOD-SV trains the FBOD model using a multi-task loss function, which consists of confidence loss and location regression loss. Specifically, during the training process, the loss value of each anchor point is obtained by calculating the weighted sum of its confidence loss and location regression loss. The formula for the loss function is as follows,
@@
$\text{L}_{\text{Total}}$ $=\frac{1}{\text{N}}\sum{\text{L}\left(\text{A}_{i}\right)}$
@@
$\text{L}_{\text{Total}}=\frac{1}{\text{N}}\left(\text{L}\left(\text{A}_{neg}\right)+\text{L}\left(\text{A}_{F_{1}}\right)+...+\text{L}\left(\text{A}_{F_{n}}\right)\right),$ (10)
@@ 3 The CPL-BC Method
The CPL-BC strategy involves maintaining two models that collaboratively select easy samples for each other, where the prediction confidence exceeds a certain threshold, and then training them. As the training progresses, the strategy gradually lowers this threshold to include more samples in the training. Specifically, the CPL-BC strategy comprises the following five steps.
@@
[…]
@@ 4 Applying the CPL-BC strategy to the training process of the FBOD model
Figure 2 illustrates the training process of the FBOD model for surveillance video using the proposed CPL-BC strategy. The process includes two main stages: the Model Prior stage and the CPL-BC training stage. In the Model Prior stage, all samples are used to train two FBOD models with randomly initialized weights, enabling them to initially acquire the ability to assess the difficulty level of samples [as shown in Figure 2(a)]. In the CPL-BC training stage, all samples are reused, and the CPL-BC strategy is applied for further training. At this stage, the models’ weights are initialized with the fixed weights obtained after the Model Prior stage [as shown in Figure 2(b)]. During this stage, the predictive confidence of one model for a sample determines the optimal sample weight for the other model’s Weighted Loss Function, controlling whether and to what extent the sample participates in training.
@@ 4.1 The Weighted Loss Function
$\text{L}_{\text{Total}}$ $=\frac{1}{\text{N}}\left(\text{L}\left(\text{A}_{neg}\right)+v_{i}\text{L}\left(\text{A}_{F_{1}}\right)+...+v_{n}\text{L}\left(\text{A}_{F_{n}}\right)\right)$
$=\frac{1}{\text{N}}\left(\text{L}\left(\text{A}_{neg}\right)+\boldsymbol{\vec{v}}\left[\text{L}\left(\text{A}_{F_{1}}\right)...\ \text{L}\left(\text{A}_{F_{n}}\right)\right]^{T}\right),$ (16)
@@ 4.3 The CPL-BC strategy for the FBOD model training
Algorithm 1 presents the pseudocode for applying the CPL-BC strategy to train the FBOD models. Initially, two FBOD models of identical structure are pre-trained using the entire dataset (model priors). During this pre-training phase, a conventional strategy is employed without differentiating sample difficulty or weighing losses. Training stops before overfitting to hard samples, typically after a relatively low number of iterations, to ensure the models initially can distinguish between easy and hard samples. Subsequently, the CPL-BC strategy is adopted for further training. This strategy aims to enable the models to progressively learn features of flying bird objects, starting with easy samples and moving to harder ones. In CPL-BC, the two models assess each other’s sample difficulty based on prediction confidence, mitigating the issue of cumulative bias that may arise when a single model selects samples.